Wildlife biologists catch, tag, and release 90 deer into a wildlife refuge. Two weeks later, they observe a sample of 154 deeer, 77 of which are tagged. Assuming the ratio of tagged deer in the sample is the same as the ratio of tagged deer in the refuge, how many total deer are in the refuge?
Answer by flame8855(424)

ratio = 77/155
let x be the number of tagged deer in the refuge
77/155 = x/90
x = 45